OpenStudy (anonymous):

Determine whether the sequence below is a geometric sequence and, if so, find a formula that describes the sequence. 1, 3, 9, 27, 81, …

OpenStudy (mathmate):

Umm... 3/1=3 9/3=3 27/9=3 81/27=3 ... Is it a geometric sequence?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

right?

OpenStudy (mathmate):

Yes, now you have to find the rule.

OpenStudy (mathmate):

knowing that the common ratio is 3.

OpenStudy (anonymous):

an= 3^n-1?

OpenStudy (anonymous):

It came out as 9.

OpenStudy (anonymous):

So its not that one.

OpenStudy (anonymous):

I dont think it is an= 3^n

OpenStudy (mathmate):

I think you had it, but the parentheses were messed up, Would you like to try again with \[a_n = 3^{n-1}\]

OpenStudy (anonymous):

an=3(1^n-1) ?

OpenStudy (anonymous):

oh ok.

OpenStudy (mathmate):

Good job!

OpenStudy (anonymous):

Thats it?

OpenStudy (mathmate):

Yep. You've got everything the question asked for. It's seems simpler than you thought, right?

OpenStudy (anonymous):

Yep. There is one more thats sorta like it that i dont get

OpenStudy (anonymous):

1, 2, 3, 4, 5, ...

OpenStudy (anonymous):

Since it is a geometric sequence, i have to multiply.

OpenStudy (mathmate):

Is it a geometric sequence?

OpenStudy (mathmate):

Or is it an arithmetic sequence?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Geometric

OpenStudy (mathmate):

Does 2/1=3/2=4/3=5/4... If not it is not a geometric sequence.

OpenStudy (anonymous):

an=2*(1)^n-1 work for it?

OpenStudy (anonymous):

Oh wait, it might not be a geometric sequence at all.

OpenStudy (mathmate):

no, it is not. It is an arithmetic sequence.

OpenStudy (anonymous):

Oh ok, thanks.

OpenStudy (mathmate):

Welcome! :)